# Information About Other Agents in Mechanism Design

### v2

@@ Abstract
We show the existence of mechanism design settings where the planner has an interest in agents receiving noisy signals about the types of other agents. When the planner is interested only in partial implementation — designing a mechanism for which there exists an equilibrium that yields a desired outcome — any social choice rule that is incentive-compatible after agents receive additional information about other agents is incentive-compatible without this information. However, additional information about other agents can eliminate undesired equilibria, making it helpful to a planner interested in full implementation — designing a mechanism for which every equilibrium outcome is desirable. We provide a sufficient condition under which a social choice rule that is not fully implementable when agents have no information about types of other agents can become fully implementable if agents have additional information.
@@ 1 Introduction
The goal of a social planner in any mechanism design problem is to design a game under which a desirable outcome emerges in equilibrium. In implementation theory, agents have private information consisting of their own preferences over a set of outcomes as well as information about the preferences of other agents. Naturally, the set of equilibrium outcomes that can emerge from a mechanism is sensitive to the details of the information available to agents. A considerable body of mechanism design literature focuses on the effects of changing information structures on achieving desirable outcomes. Most of this work focuses on how much agents know about variables that directly affect their own payoffs — for example, the information a bidder knows about her own valuation for an object (Bergemann et al. (2022)) — rather than information about other agents’ preferences.
@@
In this paper, we examine how the planner can have an interest in agents possessing additional information about the preferences of other agents, revisiting the Bayesian implementation problem as studied by Jackson (1991), as well as Postlewaite and Schmeidler (1986) and Palfrey and Srivastava (1989). In this setting, agents fully understand their own preferences over a set of outcomes but do not know the preferences of other agents. Agents’ beliefs about the preferences of other agents are drawn from a common prior, which will be assumed to have full support. This paper additionally supposes that agents can receive a full-support signal regarding the preferences of other agents. The model, which is explained in section 2, considers how the implementability of a social choice function (SCF) changes as the information structure varies. The question is whether and when there exist SCFs that are not implementable when agents have no information about the types of other agents but are implementable when one or more agents receive additional information about other agents.
@@
This question can be asked with respect to either partial or full implementation. Partial implementation requires only that the planner’s desired outcome is achieved by some equilibrium of the game, while full implementation demands that every equilibrium of the game yields a desirable outcome. It is well-known that the partial implementability of an SCF depends only on whether it satisfies incentive compatibility, which requires that it is a best response for an agent to behave truthfully if all other agents are behaving truthfully. In section 3, we show that any SCF which satisfies incentive compatibility when agents receive informative signals about the types of other agents also satisfies incentive compatibility when agents do not have this additional information. This implies that a planner concerned only with partial implementation will never prefer an agent to have additional information about other agents. The reason is when an agent has access to an informative signal about the preferences of other agents, the planner must ensure that the agent is incentivized to report truthfully in each signal realization. However, a weighted average of these constraints across signal realizations yields the incentive constraint needed for the agent if she did not have access to additional information, making this latter constraint easier to satisfy.
@@
The remainder of the paper shows how additional information about the types of other agents can help an agent signal to the planner that an undesirable equilibrium is being played, which is relevant if the planner is concerned with full implementation. In section 4 we present the main example of the paper: an SCF that is not fully implementable when agents do not have access to informative signals about the preferences of other agents but is fully implementable when agents have access to a sufficiently accurate signal about other agents. The condition that rules out undesired equilibria in Bayesian implementation settings is Bayesian monotonicity, which guarantees that agents have sufficient incentives to signal to the planner that an undesirable equilibrium strategy is being played. Much of this paper examines how Bayesian monotonicity varies with the information agents have about other agents. Section 5 presents a sufficient condition under which an SCF that is not fully implementable when agents have no information about the types of other agents becomes fully implementable under a more informative information structure.
@@
An obvious concern with endowing an agent with additional information about the preferences of other agents is that it can affect their own incentive to report truthfully to the mechanism. Indeed, the conclusion of section 3 is that incentive compatibility becomes harder to satisfy as agents receive additional information about other agents. To simplify the analysis, this paper focuses on the effects of modifying the beliefs of agents with dominant strategies. In both the main example and the general results, informative signals about other agents are only accessed by agents for which truth-telling is already a dominant strategy. This prevents the elimination of the planner’s desired equilibrium outcome after the agents receive informative signals about the types of other agents and focuses the analysis on how additional information can allow an agent to signal undesirable deviations to the planner.
@@
The signal we construct — in order to incentivize an agent $i$ to report to the planner that a specific untruthful equilibrium is being played — relies on a partition of the set of types of agents other than $i$ into two groups and informs agent $i$, with a sufficiently high degree of accuracy, whether the realized types of other agents belong to one group or the other. The intuition of this partition construction is the following: Each group should contain a type profile of other agents such that if the agent $i$ knew this type profile realized, she would prefer all other agents report truthfully rather than misreport according to the untruthful equilibrium, regardless of her own type. If agent $i$ is confident enough that the realized type profile is in one of these groups, she can propose an alternate SCF that swaps the outcomes implemented under the truthful strategy versus the untruthful strategy. By construction, this alternate SCF can only give agent $i$ a weakly lower payoff relative to the original if all other agents report truthfully, so agent $i$ would only propose it if an untruthful strategy profile is being played. Agent $i$ will have an incentive to propose this alternate SCF whenever the signal is sufficiently accurate.
Though a planner may want, for the goal of full implementation, agents to have access to informative signals about the types of other agents compared to no information, a key point of the analysis is that the planner does not always prefer a more informative signal structure. In particular, the signals constructed do not accurately inform any agent of the exact type profiles of other agents. One can imagine many reason why such a signal structure would be ineffective. For example, if there is a type profile of other agents such that all types of an agent $i$ would prefer play an untruthful equilibrium strategy as opposed to a truthful one, there will be no way to incentivize agent $i$ to report the untruthful equilibrium to the planner if she is sufficiently confident that this undesirable type profile has been realized. The conclusion that a designer may prefer information structures that are more informative, but not too much more informative, is consistent with the rapidly growing literature on information design spurred by Rayo and Segal (2010) and Kamenica and Gentzkow (2011). Thorough reviews of recent advancements in information design can be found in Bergemann and Morris (2019) and Kamenica (2019).
@@
This paper most directly contributes to the literature on full Bayesian implementation, which began with Postlewaite and Schmeidler (1986), Palfrey and Srivastava (1989), and Jackson (1991). This paper also adds to the extensive literature on the effect of changing information structures in mechanism design settings pioneered by Bergemann and Välimäki (2002). It relates especially to the literature on information disclosure in mechanism design. In auction settings, Ganuza (2004) studies the optimal way for a seller to disclose a costly public signal about a good’s characteristic in a second-price auction, which is a component of bidders’ private valuations, while Bergemann and Pesendorfer (2007) investigate the joint design problem in which the seller controls the information bidders have about their own valuations and the auction rules. Other papers that have investigated the effects of information disclosure in auction settings include Eső and Szentes (2007), and Bergemann et al. (2022). Similar questions have also been investigated in screening settings by Li and Shi (2017) and Roesler and Szentes (2017). As previously stated, much of the current literature focuses on how much agents know about their own preferences. This thesis focuses on a mechanism designer’s interest in agents knowing about other agents. Additionally, unlike the auction papers that portray the auctioneer as having the ability to provide information to agents, this thesis remains agnostic on where the information about other agents comes from. The emphasis is on how changing information structures can affect the planner’s implementation goals. Finally, while our paper asks whether a planner can be interested in agents’ information about other agents, one can also ask about an agent’s incentives to acquire such information, as done in Gleyze and Pernoud (2023), who show that most economically meaningfully mechanisms cannot prevent an agent who has the option of acquiring information about her own type as well as types of other agents from choosing to acquire information about other agents.
@@ 2 Model
The planner wishes to implement a social choice function (SCF) $F:\Theta\to A.$ The planner commits to a mechanism $\Gamma=\langle M,h\rangle$, such that $M=\prod_{i=1}^{n}M_{i}$ consists of a message space $M_{i}$ for each agent $i$ and $h:M\to A$ is the outcome function. A strategy profile $s=(s_{1},\dots,s_{n})$ specifies for each agent $i$ a function $s_{i}:\Theta_{i}\to M_{i}$. At any state $\theta$, let $s(\theta)=(s_{1}(\theta_{1}),\dots,s(\theta_{n}))$ be the message profile sent by all agents, and $s_{-i}(\theta_{-i})=(s_{j}(\theta_{j}))_{j\neq i}$ denote the messages sent by agents other than $i.$ Let $B^{\Gamma}$ be the set of Bayesian Nash equilibria for a mechanism $\Gamma.$ If there is a $s\in B^{\Gamma}$ such that $h\circ s=F$ then $\Gamma$ partially implements $F$, and $F$ is considered partially implementable. If $\Gamma$ partially implements $F$, and $B^{\Gamma}$ contains only one equilibrium outcome, then $\Gamma$ fully implements $F$ and $F$ is considered fully implementable.
@@
An SCF $F$ is incentive-compatible if and only if for all agents $i$ and all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},$
@@
where the expectation is taken over the belief of type $\theta_{i}$ about $\Theta_{-i}.$
@@
It was shown in Jackson (1991) that incentive compatibility and Bayesian monotonicity are necessary and sufficient for an SCF $F$ to be fully implementable in economic environments with at least 3 agents. [^2] An environment is economic if given any state $\theta\in\Theta$ and SCF $F$, at least two agents prefer some outcome $a\in A$ to $F(\theta).$ [^3] Jackson (1991) also identifies a closure condition for collections of SCFs, social choice sets, but the condition is vacuous for the case of a single-valued SCF. While incentive compatibility guarantees that truth-telling is an equilibrium of the direct mechanism, Bayesian monotonicity is the condition that allows the planner to design a mechanism that can eliminate undesired equilibria. The key intuition behind this condition is as follows. Suppose agents are initially at an equilibrium that uses a deception $\alpha$ such that $F\circ\alpha\neq F.$ By allowing any individual agent $i$ to specify an alternate SCF (the function $y$ can be extended to $\Theta$ by implementing the same outcome regardless of agent $i$’s type) that is implemented only if it appears to harm agent $i$, an agent can choose to use this alternate SCF to signal to the planner that the agents are lying. Upon seeing the alternate SCF, the planner understands that a deception is being played because it would harm agent $i$ in a truth-telling equilibrium, regardless of her own type. Bayesian monotonicity gives at least one type of agent $i$ an incentive to specify an alternate SCF because she would prefer the alternate SCF to $F$ under the deception $\alpha$, ruling out the deception as an equilibrium. Bayesian monotonicity guarantees that such a procedure can be done for every undesirable deception.
@@
where $\zeta_{i}$ is a finite set of signal realizations and $\sigma_{i}(\theta_{-i})$ is the signal observed by agent $i$ when the type profile of other agents is $\theta_{-i}.$ The notation $\Delta(\zeta_{i})$ denotes probability distributions over $\zeta_{i}$. Assume that $\sigma_{i}$ has full support over $\zeta_{i}$ for every agent $i.$ After observing the realization of $\sigma_{i}$ and her type $\theta_{i}$, agent $i$ updates her beliefs regarding $\Theta_{-i}$ (and $\Theta$) according to Bayes’ rule. In particular, suppose $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ and agent $i$ has type $\theta_{i}.$ Upon observing $\sigma_{i}=A_{ij}$, agent $i$ places probability
@@
on type $\theta_{-i}.$ Note that no signal $\sigma_{i}$ can be conditioned on agent $i$’s type.
@@
We make several substantive assumptions that simplify and clarify the analysis. Other than requiring all $\sigma_{i}$ to have full support, we assume there are no restrictions over the structure of the signal agent $i$ can receive. Thus, agents do not have private information that can constrain the types of signals they can process. Additionally, we assume the signals are not costly in order to focus on the role of the additional information regarding other agents’ types. Finally, the structure of the signals $\{\sigma_{i}\}_{i\in I}$ is assumed to be common knowledge.
@@ 2.1 The full-support signal assumption
To justify the assumption that each $\sigma_{i}$ has full support, it is helpful to consider the following example. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There is a uniform common prior, so all states are equally likely. There are 5 possible outcomes $A=\{a,b,c,d,e\}.$ The SCF $F$ is as follows.
@@
If $\sigma_{i}$ is uninformative for all $i$, it follows that $F$ is not fully or partially implementable under the specified information structure. On the other hand, if $\sigma_{i}$ is perfectly informative for all $i$, each agent $i$ has complete information regarding the state $\theta$ upon observing the realization of $\sigma_{i}$ and learning their own type $\theta_{i}$. In this case, the problem becomes the classical Nash implementation problem solved by Maskin (1999), which shows that Maskin monotonicity [^5] An SCF $F$ is Maskin-monotonic if and only if for all $\theta,\theta^{\prime}\in\Theta$, whenever $F(\theta)\neq F(\theta^{\prime})$, there exist $i\in I$ and $a\in A$ such that $u_{i}(F(\theta),\theta)\geq u_{i}(a,\theta)$ but $u_{i}(F(\theta),\theta^{\prime})<u_{i}(a,\theta^{\prime})$. and no veto power [^6] An SCF $F$ satisfies no veto power if for all $(a,j,\theta)\in A\times I\times\Theta$, the condition $u_{i}(a,\theta)\geq u_{i}(b,\theta)$ for all $b\in A$ and $i\neq j$ implies $a=F(\theta).$ are sufficient conditions for the full implementability of $F$ when there are at least three agents. No veto power is automatically satisfied because $F$ never allows two agents to simultaneously get their most preferred outcome. To see that $F$ is Maskin-monotonic, first note that outcomes $a$ and $c$ fall relative to each other in agent 3’s preference ranking when moving between types $\theta_{3}$ and $\theta_{3}^{\prime}$ and vice versa. Additionally, $b$ falls relative to $a$ in agent 1’s preference ranking when moving from type $\theta_{1}^{\prime}$ to $\theta_{1}$, and $a$ and $c$ both fall relative to $d$ when agent 1 moves from type $\theta_{1}$ to $\theta_{1}^{\prime}.$ Finally, $a$ and $c$ fall relative to each other in agent 2’s preference ranking when moving between types $\theta_{2}$ and $\theta_{2}^{\prime}$ and vice versa, and $b$ falls relative to $a$ and $c$ when moving from $\theta_{2}^{\prime}$ to $\theta_{2}.$ Thus $F$ is implementable if each $\sigma_{i}$ provides agent $i$ with a perfect signal regarding $\Theta_{-i}.$
@@
This example demonstrates that agents receiving information about the preferences of others can allow an initially non-implementable SCF to become implementable, a possibility that has not yet been noted in the literature. It also shows how the incentive compatibility condition changes when agents are able to completely rule out certain states of the world. In particular, since implementation with complete information is a special case of Bayesian implementation, it must be that $F$ satisfies incentive compatibility under complete information even though it is not incentive-compatible under the uniform common prior. This is because an agent’s type $\theta_{i}$ encapsulates information about both the preferences of agent $i$ and the information she has about the types of other agents. In the example above, $\theta_{i}$ is a complete description of agent $i$’s type under the uniform common prior, because it specifies her preferences over the outcomes but she is not able to rule out any type profile of the other agents. Under complete information, however, $\theta_{i}$ is no longer an adequate description of her type. Since agent $i$ is able to completely discern the types of the other agents, her type includes what she knows to be the type profile of the other agents. As such, if one were to translate the example above with complete information into the Bayesian implementation setup, then $\Theta_{i}=\prod_{j=1}^{3}\{\theta_{j},\theta_{j}^{\prime}\}$ for all $i.$ So, when agent $i$ has preferences according to $\theta_{i},$ she has four corresponding types $t_{i}\in\Theta_{i}$. This is consistent with the mechanism exhibited in Maskin (1999), where each agent was asked to report the complete state of the world, which included information about the preferences of all agents.
@@
In the translation of a complete information implementation problem into one of Bayesian implementation, the states $\Theta=\prod_{i}\Theta_{i}$ contain many zero-probability states. In fact, any $t=(t_{1},t_{2},t_{3})\in\Theta$ for which each $t_{1}=t_{2}=t_{3}$ does not hold is a state that occurs with zero probability. However, the planner still specifies the SCF for these events in order to preclude equilibria that are deceptions. This renders the incentive compatibility condition trivial when $n\geq 3$, because if one agent unilaterally deviates from the truth, the planner can simply ignore the deviation and implement the outcome in the state reported by $n-1$ agents.  [^7] This opens up the possibility of undesired equilibria, which are ruled out by Maskin monotonicity. The Bayesian monotonicity condition reduces to Maskin monotonicity in complete information environments. As such, the restriction that each $\sigma_{i}$ always has full support simply prevents changes to any type space $\Theta_{i}$ after agents receive the signal realization.

### v3

@@ Abstract
We study mechanism design settings where the planner has an interest in agents receiving noisy signals about the types of other agents. When the planner is interested only in partial implementation — designing a mechanism for which there exists an equilibrium that yields a desired outcome — any social choice rule that is incentive-compatible after agents receive additional information about other agents is incentive-compatible without this information. However, additional information about other agents can eliminate undesired equilibria, making it helpful to a planner interested in full implementation — designing a mechanism for which every equilibrium outcome is desirable. We provide a sufficient condition under which a social choice function that is not fully implementable when agents have no information about types of other agents can become fully implementable if agents have additional information.
@@ 1 Introduction
The goal of a social planner in any mechanism design problem is to design a game under which a desirable outcome emerges in equilibrium. Agents have private information about their own preferences and preferences of other agents. Naturally, the set of equilibrium outcomes that can emerge from a mechanism is sensitive to the details of the information available to agents. A considerable body of mechanism design literature focuses on the effects of changing information structures on achieving desirable outcomes. Most of this work focuses on how much agents know about variables that directly affect their own payoffs — for example, the information a bidder knows about her own valuation for an object (Bergemann et al. (2022)) — rather than information about other agents’ preferences.
@@
In this paper, we examine how the planner can have an interest in agents possessing additional information about the preferences of other agents, revisiting the Bayesian implementation problem as studied by Jackson (1991) (see also Postlewaite and Schmeidler (1986) and Palfrey and Srivastava (1989)). In this setting, agents fully understand their own preferences over a set of outcomes but do not know the preferences of other agents. We consider how the implementability of a social choice function (SCF) changes if agents can receive a full-support signal regarding the preferences of other agents. In particular, our main result gives a sufficient condition under which a planner interested in full implementation — designing a mechanism such that every equilibrium yields a desirable outcome — will strictly prefer an agent to have additional information about agents.
@@
In section 3, we show that any SCF which satisfies incentive compatibility when agents receive informative signals about the types of other agents also satisfies incentive compatibility when agents do not have this additional information. This implies that a planner concerned only with partial implementation — ensuring that the desirable outcome constitutes an equilibrium of the mechanism — will never strictly prefer an agent to have additional information about other agents. The argument is straightforward: When an agent has access to an informative signal about the preferences of other agents, the planner must ensure that the agent is incentivized to report truthfully in each signal realization. A weighted average of these constraints across signal realizations yields the incentive constraint needed for the agent if she did not have access to additional information, making this latter constraint easier to satisfy.
@@
The remainder of the paper examines how additional information about the types of other agents can help an agent signal untruthful equilibria to the planner. In section 4, we present the main example of the paper: an SCF that is not fully implementable when agents have no information about the preferences of other agents but is fully implementable when agents have more information. The condition that rules out undesired equilibria in Bayesian implementation settings is Bayesian monotonicity, which guarantees that agents have sufficient incentives to signal to the planner that an undesirable equilibrium strategy is being played. We focus heavily on how Bayesian monotonicity varies with the information agents have about other agents. Section 5 presents a sufficient condition for an SCF that is not fully implementable when agents have no information about the types of other agents to be fully implementable when agents have more information.
@@
An obvious concern with giving an agent additional information about types of other agents is that it can affect their own incentive to report truthfully to the mechanism. Indeed, the conclusion of section 3 is that incentive compatibility becomes (weakly) harder to satisfy if agents receive additional information about other agents. To avoid eliminating truthful reporting as an equilibrium, we only modify the information available to agents for which truth-telling is a dominant strategy in the direct mechanism of a given SCF. Our results do not speak to SCFs for which no agent has a dominant strategy in the direct mechanism.
@@
To establish our main result, we note that an incentive-compatible SCF is not fully implementable if and only if there is an untruthful equilibrium of the direct mechanism such that no agent has an incentive to report the deception to the planner. The task is then to construct a signal that gives some agent $i$ the ability to eliminate these equilibria. In general, if truth-telling is a dominant strategy for player $i$, an untruthful equilibrium $\alpha$ of the direct mechanism will involve some subset of other agents playing deceptions. The signal we construct partitions the set of other agents’ types into two groups. Each group in the partition must contain a type profile of other agents such that if the agent $i$ knew this type profile realized, she would prefer all other agents to report truthfully rather than deceive according to $\alpha$, regardless of her own type. If agent $i$ is confident enough that the realized type profile is in one group or the other, she can propose an alternate SCF that swaps the outcomes implemented under the truthful strategy versus the deception. By construction, this alternate SCF can only give agent $i$ a weakly lower payoff relative to the original if all other agents report truthfully, so agent $i$ would only propose it if $\alpha$ is being played. Whenever the signal is sufficiently accurate, agent $i$ can eliminate $\alpha$ as an equilibrium. We conclude by showing that this procedure can iteratively eliminate all untruthful equilibria of the direct mechanism.
Though a planner may want agents to have access to informative signals about the types of other agents compared to no information, our analysis also demonstrates that a planner may prefer an information structure that lies strictly between no information and complete information. This conclusion is consistent with the rapidly growing literature on information design spurred by Rayo and Segal (2010) and Kamenica and Gentzkow (2011).
@@
This paper directly contributes to the literature on full Bayesian implementation — which began with Postlewaite and Schmeidler (1986), Palfrey and Srivastava (1989), and Jackson (1991) — by shedding light on the relationship between Bayesian monotonicity and the information available to agents. Since agents are receiving informative signals about the state of the world, we also add to the extensive literature on the effects of changing information structures in mechanism design settings pioneered by Bergemann and Välimäki (2002). It relates especially to the literature on information disclosure in mechanism design. In auction settings, Ganuza (2004) studies the optimal way for a seller to disclose a costly public signal about a good’s characteristic in a second-price auction, which is a component of bidders’ private valuations, while Bergemann and Pesendorfer (2007) investigate the joint design problem in which the seller controls the information bidders have about their own valuations and the auction rules. Other papers that have investigated the effects of information disclosure in auction settings include Eső and Szentes (2007), and Bergemann et al. (2022). Similar questions have also been investigated in screening settings by Li and Shi (2017) and Roesler and Szentes (2017). Since much of the existing literature studies cases when agents’ types are independently drawn, existing results focus on how much agents know about their own preferences. This paper focuses on a mechanism designer’s interest in agents’ information about types of other agents. Additionally, unlike the auction papers that portray the auctioneer as having the ability to provide information to agents, this paper remains agnostic on where the information about other agents comes from. The emphasis is on how changing information structures can affect the planner’s implementation goals. Finally, while our paper asks whether a planner can be interested in agents’ information about other agents, one can also ask about an agent’s incentives to acquire such information, as done in Gleyze and Pernoud (2023), who show that most mechanisms incentivize agents to acquire information about other agents in environments where agents have uncertainty about their own types as well as types of other agents.
@@ 2 Model
The planner wishes to implement a social choice function (SCF) $F:\Theta\to A.$ The planner commits to a mechanism $\Gamma=\langle M,h\rangle$, such that $M=\prod_{i=1}^{n}M_{i}$ consists of a message space $M_{i}$ for each agent $i$ and $h:M\to A$ is the outcome function. A strategy profile $s=(s_{1},\dots,s_{n})$ specifies for each agent $i$ a function $s_{i}:\Theta_{i}\to\Delta(M_{i})$. At any state $\theta$, let $s(\theta)=(s_{1}(\theta_{1}),\dots,s(\theta_{n}))$ be the message profile sent by all agents, and $s_{-i}(\theta_{-i})=(s_{j}(\theta_{j}))_{j\neq i}$ denote the messages sent by agents other than $i.$ Let $B^{\Gamma}$ be the set of Bayesian Nash equilibria for a mechanism $\Gamma.$ If there is a $s\in B^{\Gamma}$ such that $h\circ s=F$ then $\Gamma$ partially implements $F$, and $F$ is considered partially implementable. If $\Gamma$ partially implements $F$, and $B^{\Gamma}$ contains only one equilibrium outcome, then $\Gamma$ fully implements $F$ and $F$ is considered fully implementable.
@@
An SCF $F$ is (Bayesian) incentive-compatible if and only if for all agents $i$ and all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},$
@@
where the expectation is taken over the belief of type $\theta_{i}$ about $\Theta_{-i},$ which is formed from the common prior $\mu$.
@@
It was shown in Jackson (1991) that incentive compatibility and Bayesian monotonicity are necessary and sufficient for an SCF $F$ to be fully implementable in economic environments with at least 3 agents. [^2] An environment is economic if given any state $\theta\in\Theta$ and SCF $F$, at least two agents prefer some outcome $a\in A$ to $F(\theta).$ [^3] Jackson (1991) also identifies a closure condition for collections of SCFs, social choice sets, but the condition is vacuous for the case of a single-valued SCF. While incentive compatibility guarantees that truth-telling is an equilibrium of the direct mechanism, Bayesian monotonicity is the condition that allows the planner to design a mechanism that can eliminate undesired equilibria. The intuition behind the condition, which we will reference throughout this paper, is as follows. Suppose agents are initially at an equilibrium that uses a deception $\alpha$ such that $F\circ\alpha\neq F.$ By allowing any individual agent $i$ to specify an alternate SCF (the function $y$ can be extended to $\Theta$ by implementing the same outcome regardless of agent $i$’s type) that is implemented only if it appears to harm agent $i$, an agent can choose to use this alternate SCF to signal to the planner that the agents are lying. Upon seeing the alternate SCF, the planner understands that a deception is being played because it would harm agent $i$ in a truth-telling equilibrium, regardless of her own type. Bayesian monotonicity gives at least one type of agent $i$ an incentive to specify an alternate SCF because she would prefer the alternate SCF to $F$ under the deception $\alpha$, ruling out the deception as an equilibrium. Bayesian monotonicity guarantees that such a procedure can be done for every undesirable deception.
@@
where $\zeta_{i}$ is a finite set of signal realizations and $\sigma_{i}(\theta_{-i})$ is the signal observed by agent $i$ when the type profile of other agents is $\theta_{-i}.$ Assume that $\sigma_{i}$ has full support over $\zeta_{i}$ for every agent $i.$ After observing the realization of $\sigma_{i}$ and her type $\theta_{i}$, agent $i$ updates her beliefs regarding $\Theta_{-i}$ (and $\Theta$) according to Bayes’ rule. In particular, suppose $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ and agent $i$ has type $\theta_{i}.$ Upon observing $\sigma_{i}=A_{ij}$, agent $i$ places probability
@@
on type $\theta_{-i}.$
@@
Our setup makes several substantive assumptions. Other than requiring all $\sigma_{i}$ to have full support, we assume there are no restrictions over the structure of the signal agent $i$ can receive. Thus, agents do not have private information that can constrain the types of signals they can process. Additionally, we assume the signals are not costly in order to focus on the role of the additional information regarding other agents’ types. Finally, the structure of the signals $\{\sigma_{i}\}_{i\in I}$ is assumed to be common knowledge.
@@ 2.1 The full-support signal assumption
The restriction that each $\sigma_{i}$ has full support is a simplifying assumption that avoid complications to the type space $\Theta$ after agents receive their signals. To illustrate we make this assumption, consider the following example under which agents receive perfectly informative signals. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There is a uniform common prior, so all states are equally likely. There are 5 possible outcomes $A=\{a,b,c,d,e\}.$ The SCF $F$ is as follows.
@@
If $\sigma_{i}$ is uninformative for all $i$, it follows that $F$ is not fully or partially implementable under the specified information structure. On the other hand, if $\sigma_{i}$ is perfectly informative for all $i$, each agent $i$ has complete information regarding the state $\theta$ upon observing the realization of $\sigma_{i}$ and learning their own type $\theta_{i}$. Then incentive compatibility becomes trivial when $n\geq 3$. Each agent’s private information now consists of the complete state $\theta$. Thus if one agent unilaterally deviates from reporting their private information truthfully, the planner can simply ignore the deviation and implement the outcome in the state reported by $n-1$ agents. In this case, full implementability of $F$ reduces to the classical Nash implementation problem solved by Maskin (1999), which proves that Maskin monotonicity [^5] An SCF $F$ is Maskin-monotonic if and only if for all $\theta,\theta^{\prime}\in\Theta$, whenever $F(\theta)\neq F(\theta^{\prime})$, there exist $i\in I$ and $a\in A$ such that $u_{i}(F(\theta),\theta)\geq u_{i}(a,\theta)$ but $u_{i}(F(\theta),\theta^{\prime})<u_{i}(a,\theta^{\prime})$. and no veto power [^6] An SCF $F$ satisfies no veto power if for all $(a,j,\theta)\in A\times I\times\Theta$, the condition $u_{i}(a,\theta)\geq u_{i}(b,\theta)$ for all $b\in A$ and $i\neq j$ implies $a=F(\theta).$ are sufficient conditions for the full implementability of $F$ when there are at least three agents. No veto power is automatically satisfied because $F$ never allows two agents to simultaneously get their most preferred outcome. To see that $F$ is Maskin-monotonic, first note that outcomes $a$ and $c$ fall relative to each other in agent 3’s preference ranking when moving between types $\theta_{3}$ and $\theta_{3}^{\prime}$ and vice versa. Additionally, $b$ falls relative to $a$ in agent 1’s preference ranking when moving from type $\theta_{1}^{\prime}$ to $\theta_{1}$, and $a$ and $c$ both fall relative to $d$ when agent 1 moves from type $\theta_{1}$ to $\theta_{1}^{\prime}.$ Finally, $a$ and $c$ fall relative to each other in agent 2’s preference ranking when moving between types $\theta_{2}$ and $\theta_{2}^{\prime}$ and vice versa, and $b$ falls relative to $a$ and $c$ when moving from $\theta_{2}^{\prime}$ to $\theta_{2}.$ Thus $F$ is implementable if each $\sigma_{i}$ provides agent $i$ with a perfect signal regarding $\Theta_{-i}.$
@@
This example demonstrates that agents receiving information about the preferences of others can allow an initially non-implementable SCF to become implementable. It also shows how the incentive compatibility condition changes when agents are able to completely rule out certain states of the world. In particular, since implementation with complete information is a special case of Bayesian implementation, it must be that $F$ satisfies incentive compatibility under complete information even though it is not incentive-compatible under the uniform common prior. This is because an agent’s type $\theta_{i}$ encapsulates information about both the preferences of agent $i$ and the information she has about the types of other agents. In the example above, $\theta_{i}$ is a complete description of agent $i$’s type under the uniform common prior, because it specifies her preferences over the outcomes but she is not able to rule out any type profile of the other agents. Under complete information, however, $\theta_{i}$ is no longer an adequate description of her type. Since agent $i$ is able to completely discern the types of the other agents, her type includes what she knows to be the type profile of the other agents. As such, if one were to translate the example above with complete information into the Bayesian implementation setup, then $\Theta_{i}=\prod_{j=1}^{3}\{\theta_{j},\theta_{j}^{\prime}\}$ for all $i.$ So, when agent $i$ has preferences according to $\theta_{i},$ she has four corresponding types $t_{i}\in\Theta_{i}$. This is consistent with the mechanism exhibited in Maskin (1999), where each agent was asked to report the complete state of the world, which included information about the preferences of all agents.
@@
In the translation of a complete information implementation problem into one of Bayesian implementation, the states $\Theta=\prod_{i}\Theta_{i}$ contain many zero-probability states. In fact, any $t=(t_{1},t_{2},t_{3})\in\Theta$ for which each $t_{1}=t_{2}=t_{3}$ does not hold is a state that occurs with zero probability. However, the planner still specifies the SCF for these events in order to preclude equilibria that are deceptions, which trivializes incentive-compatibility when $n\geq 3$. As such, the restriction that each $\sigma_{i}$ always has full support simply prevents changes to any type space $\Theta_{i}$ after agents receive the signal realization.